### v1

# Integrated Communication, Navigation, and Remote Sensing in LEO Networks with Vehicular Applications

Abstract
Traditionally, communication, navigation, and remote sensing (CNR) satellites are separately performed, leading to resource waste, information isolation, and independent optimization for each functionality. Taking future automated driving as an example, it faces great challenges in providing high-reliable and low-latency lane-level positioning, decimeter-level transportation observation, and huge traffic sensing information downloading. To this end, this article proposes an integrated CNR (ICNR) framework based on low earth orbit (LEO) satellite mega-constellations from the perspective of vehicular applications. After introducing the main working principles of the CNR functionalities to serve as the technological basis, we characterize the potentials of the integration gain in vehicular use cases. Then, we investigate the ICNR framework in different integration levels, which sheds strong light on qualitative performance improvement by sophisticatedly sharing orbit constellation, wireless resource, and data information towards meeting the requirements of vehicular applications. We also instantiate a fundamental numerical case study to demonstrate the integration gain and highlight the main tradeoffs in managing the ICNR networks from the perspective of vehicular applications.
Index Terms:
Low earth orbit, mega-constellation, navigation, remote sensing, vehicular applications.
I Introduction
In the past thirty years, the ever growing human demand on seamless information coverage, terminal positioning, and earth observation makes global artificial satellite industry evolve and expand significantly, leading to three main independent bodies: communication, navigation, and remote sensing (CNR) satellite systems. Due to different focuses, the existing satellites are operating in various orbits, including geo-stationary earth orbit (GEO) with an altitude of 35,786 km, medium earth orbit (MEO) with an altitude of from 2,000 km to 20,000 km, and low earth orbit (LEO) with an altitude of less than 2,000 km. Specifically, the UK OneWeb communication systems, the USA global positioning system (GPS), and China Gaofen 4 remote sensing satellites are working in LEO, MEO, GEO orbits, respectively [1].
While deciding the altitude of a single satellite, there is a fundamental tradeoff between coverage range and signal propagation latency. However, by developing a huge constellation of low-altitude satellites, the coverage can be enhanced while reducing latency, which brings about a rising tide of world-wide LEO satellite network construction in the past ten years. Meanwhile, technological advancements have made it possible to build small cost-effective LEO satellites, develop reusable rockets, and adjust moving status in orbit. In this context, the economically feasible LEO satellite manufacture, launching, and in-orbit management techniques have enabled many countries, organizations, and commercial companies to establish their own satellite constellations. Besides the advantage of low latency, the huge constellation of LEO satellites is also good at enhancing capacity and reliability of communication, accuracy and rapidity of positioning, and spatial and temporal resolution of remote sensing [2]. Due to such a great potential, the number of LEO satellites may persistently grow towards the capacity of the orbits, predicted as several millions, leading to mega-constellations of LEO satellites in the future [3].
It had long been the situation that different kinds of satellites usually serve the customers with distinct purposes. Particularly, communication satellites mainly provide special users connectivity in regions out of terrestrial network coverage, such as sea, forest, and desert areas. Navigation satellites offer global positioning, navigation, and timing (PNT) services to a large body of location-dependent applications, such as object tracking, route planning, and industrial automation. Remote sensing satellites are able to provide a top view of the earth, which can be applied to agriculture monitoring, natural disaster discovering, and traffic network optimization. However, heterogeneous services are required by vehicular terminals, especially in the era of autonomous driving. For example, an automated connected vehicle in intelligent transportation systems may require timely remote-sensing traffic information to optimize its route planning, necessitate high-precision positioning to perform lane-level navigation, and need high-capacity communication to transfer diversified messages including remote sensing data [4]. To provide the required CNR services to future vehicular applications, the LEO mega-constellation of satellites is expected to be a promising solution.
Figure 1: Fundamental working principles of satellite communication, navigation, and remote sensing.
Traditional separate-deployed CNR satellite systems face considerable challenges in meeting captious vehicular requirements due to loss of sharing gain in terms of orbit constellation, wireless resource, and data information, which further prohibits joint CNR optimization towards satisfying the autonomous driving application. In this regard, the framework of integrated communication and navigation has been proposed. In the early stage of such integration dating back to 1980s, high-altitude satellites are mainly designed for vehicular navigation, implying that navigation is taken as the primary function and communication is introduced as the secondary capability. With the proliferation of LEO satellite based broadband Internet, the integration moves the foundation to the communication side, where navigation data is incorporated into the communication packets [5]. Another topic is the integration of communication and remote sensing, which has hardly been considered until the recent flourishing low-altitude space networks, mainly because the delay requirement of remote sensing data downloading for traditional users is not as strict as present applications, such as intelligent transportation. Particularly, communication among remote sensing satellites offers the opportunity of onboard sensing information fusion, leading to higher resolution and quicker response, which is pretty beneficial to modern vehicular users [6]. Other than the combination of dual functionalities, a few works have begun to consider the combination of the three functionalities. Frequency diverse array and orthogonal frequency division multiplexing have been utilized in [7] to efficiently exploit waveform diversity for integrated communication, sensing, and navigation. In [8], the constellation design has been investigated and a simplified analysis has been provided to jointly consider the communication and remote sensing requirements. The attractive prospect of the integration with applications in ecological environment monitoring has been demonstrated in [9]. Last but not least, the national aeronautics and space administration (NASA) has also raised its interests on multi-purpose space cubes with remote sensing, communication, and navigation integration [10].
Despite sporadically mentioned in the literature, the integration of the CNR in a mega-constellation of LEO satellite networks has not been completely investigated with its fundamentals, classification, potentials, and challenges being specified. In this article, we first summarize the most fundamental principles of the CNR techniques individually, followed by demonstration of the proposed integrated CNR (ICNR) framework with vehicular applications as a representative use case. Then, different levels of integration in the proposed ICNR network are discussed and a fundamental numerical case study is conducted to demonstrate the integration gain. Finally, we highlight several main tradeoffs that need to be elaborated in practical ICNR network deployment and management.
II ICNR Framework with Vehicular Applications
In this section, we first present the fundamental principles of the satellite CNR systems, as described in Fig. 1. Then, an ICNR framework is proposed with typical use cases illustrated in vehicular applications.
II-A Satellite Communication
Traditional communication satellites usually act as relays to transit information, i.e., forwarding information from a source node to a destination node on the earth, where the destination node can be as far as tens of thousands of kilometers away from the source node. Taking downlink transmission as an example, we consider terminal 0 as the reference user receiving information signal from satellite 0. Without loss of generality, there are additional $M$ satellites in the zenith of the terminal, causing interference to the terminal. Denote $x_{m}(t)$ as the transmitted signal from satellite $m$ at time $t$. Then, the received signal at time $t$ can be expressed as
$y_{0}(t)=h_{0}x_{0}(t-\tau_{0})+\sum_{m=1}^{M}h_{m}x_{m}(t-\tau_{m})+n(t),$ (1)
where $n(t)$ is the receiving noise, and $\tau_{m}$ and $h_{m}$ are the propagation delay and channel coefficient from satellite $m$ to the terminal, respectively. The main task of satellite communication is to recover $x_{0}(t)$ from $y_{0}(t)$ by removing the interference and noise, i.e., the second and third terms in (1), respectively. Generally, capacity, reliability, and latency are the three dominant metrics that are usually cared about by vehicular users, e.g., to download the remote sensing data. A peculiarity of vehicular applications is the mobility of terminals, which makes channel features to be different from static users and thus deserves special attention in practical systems.
II-B Satellite Navigation
Positioning and timing are the two main tasks of navigation systems, which are the most fundamental requirements of vehicular terminals. Since the ephemeris and almanac of satellites are available and atomic clocks are equipped in satellites, navigation satellites can be aware of their positions as well as time. Consider that a target user aims to determine the time $t$ and its position denoted by three-dimensional coordinates $(u,v,w)$ in a Cartesian coordinate system. To this end, upon the user receiving a signal from satellite $k$ emitted at time $t_{k}$, it is able to get the distance from itself to satellite $k$ being $c(t-t_{k})$, where $c$ is the speed of light and $t_{k}$ is the time at which the signal is emitted by satellite $k$. Using simple trigonometry, we have the equation given by
$(u-u_{k})^{2}+(v-v_{k})^{2}+(w-w_{k})^{2}=c^{2}(t-t_{k})^{2},$ (2)
where $(u_{k},v_{k},w_{k})$ is the position of satellite $k$. Since the user needs to determine four unknown parameters, $u$, $v$, $w$, and $t$, at least four satellites are required to establish four equations as in (2) for $k=1,2,3,4$. Due to ranging errors caused by inaccurate reference satellite position, multi-path propagation, light refraction, etc., the position and time are estimated rather than computed exactly. Generally, receiving signals from additional satellites does positioning and timing a favor with more precision by finding a least-square solution of (2) with more than four available navigation satellites. Besides the number of visible satellites, the relative geometric relationship among user equipment and visible satellites also plays a significant role in determining the accuracy of positioning and timing, usually measured by the geometric dilution of precision (GDOP) indicating the ratio of positioning and timing error to ranging error [11]. Vehicular terminals can also leverage inertial measurement units (IMU) and cooperative technique to improve the positioning accuracy in addition to using the available satellite navigation signals.
Figure 2: Illustration of the proposed ICNR framework with vehicular applications considered in representative rural and urban use cases.
II-C Satellite Remote Sensing
Remote sensing satellites are designed to view the earth from a position in orbit and usually produce images of the earth surface. These images are crucial to a diverse array of applications, including environmental monitoring, weather prediction, agriculture management, disaster detection, and intelligent transportation. There are generally two kinds of satellite remote sensing technologies, namely passive and active ones. The passive remote sensing satellites detect energy emitted or reflected from objects on the earth, usually operating in visible, infrared, thermal infrared, and microwave portions of the electromagnetic spectrum. Since the passive sensing technology is vulnerable to lighting conditions and dependent on ambient energy, the active remote sensing, especially synthetic aperture radar (SAR) based remote sensing, plays an important role in mission-critical applications. Similar to general radar remote sensing mechanism, a SAR satellite performs electromagnetic radiations to illuminate the target on the earth and subsequently records the reflected radiation signal to get earth observations. This implies that it is difficult for a remote sensing satellite to achieve high-resolution images by very narrow beam focus due to antenna size and working frequency limitations. However, by exploiting the motion of spacecrafts, SAR satellites can achieve equivalent long-aperture antennas by accumulating signal reception over time, leading to an azimuth resolution $\delta_{\rm a}=\frac{d_{\rm a}}{2}$, where $d_{a}$ is the aperture length of the antenna in the azimuth direction. Although a SAR satellite prefers a small $d_{a}$, practical antenna size should be large enough to restrict the illuminated ground area to avoid ambiguous echo signal reception [12]. With the pulse compression technique, the cross-track resolution of a SAR satellite can be estimated as
$\delta_{\rm c}=\frac{c}{2B\sin\varphi},$ (3)
where $B$ is the spectrum bandwidth and $\varphi$ is the view angle, i.e., the satellite centered angle between the ground target and the nadir. In vehicular use cases, the spatial resolution should be low enough to distinguish different vehicles. Besides the spatial precision, the temporal precision also deserves special attention, which is traditionally captured by the revisit time. It is worth noting that the temporal resolution should be defined with respect to the whole system in a mega-constellation of LEO satellites, i.e., the time it takes for a target to be observed twice by the constellation rather than by a satellite. The temporal resolution needs to be as small as the variation period of the road traffic in vehicular applications, which can be measured by age of information (AoI) of the remote sensing data received by vehicles [13].
Figure 3: Illustration of the different integration levels of the proposed ICNR framework.
II-D Proposed ICNR Framework
Since all of the CNR satellites involve electromagnetic signal transmitting and receiving, we propose a LEO satellite based ICNR framework by integrating the three distinct functionalities in a single satellite and constructing an ICNR network with inter-satellite links (ISLs). Such an ICNR system is operationally feasible thanks to ever developing satellite manufacturing and launching technologies. Actually, the Iridium-next system is planning to embrace communication payload, navigation enhancement payload, and earth observation sensors within a single satellite [14]. More importantly, orbit, spectrum, time, and energy resource can be more efficiently exploited compared with the traditional physically separate and independent systems.
Particularly, operating in a mega-constellation LEO satellite networks, the ICNR framework sheds new light on rural and urban vehicular applications, as illustrated in Fig. 2. In case of being trapped on roads due to natural diaster in distant area without terrestrial infrastructure coverage, the vehicle can first get its enhanced localization by exploiting more visible LEO satellites and then send an save-our-ship (SOS) signal to an available satellite, which further forwards the SOS signal to a ground rescue station. On the other hand, an appropriate satellite can be triggered to perform real-time remote sensing for the related region and then send the disaster sensing information to the ground rescue station. Based on the geographical disaster condition and the location of the stranded vehicle, the best rescue plan can be made, e.g., determining the best transportation means and the rescue trajectories.
The ICNR networks are also quite beneficial to future intelligent transportation systems (ITS) even within the coverage of terrestrial networks, as presented in the urban area in Fig. 2. Because of stochastic traffic distribution, the shortest driving path may not be the best choice for a vehicle. In this context, a satellite may first implement remote traffic sensing for the area covering both the current location and the destination point. Then, such sensed traffic information is forwarded through ISLs to other satellites, which are closer to the vehicle or the ITS scheduling center. After that, a proper relaying satellite can transfer the traffic distribution information to the vehicle to let it make a better route plan. Alternatively, a relaying satellite may also send the traffic distribution information to the ITS scheduling center, which produces a route suggestion after considering the whole traffic system and notifies the vehicle. On the other hand, it becomes feasible to achieve more enhanced centimeter-level positioning by exploiting more visible LEO satellites and simultaneous multi-purpose signal transmission. Finally, based on the timely traffic congestion information, the accurate positioning, and the suggestion from the ITS center, the vehicle is able to make the optimal route decision and lane-level navigation. In summary, real-time vehicle-distinguishable remote sensing, lane-level positioning, and huge remote sensing information downloading are commonly required by intelligent connected vehicles, which may be suitably addressed by the proposed ICNR networks with a mega-constellation of LEO satellites.
III Integration Levels of ICNR
Depending on the coupling degrees among different functionalities, we classify the proposed ICNR framework into four stages, i.e., function-level, signal-level, data-level, and application-level integration, in an increasing order of the coupling degree, as summarized in Fig. 3.
  • L1: Function-Level Integration
Function integration stands for the lowest integration level of the ICNR networks, where the CNR functionalities are loaded in a single LEO satellite but operate almost independently. Each functionality uses its exclusive hardware and software components to accomplish its own task, except some underlying equipment, such as the battery system and central on-board processing unit that are shared by the CNR. This means that the ICNR networks can be sometimes treated as three mutually independent systems for the CNR, with a common constellation shared. As the equivalent number of satellites for each functionality increases by such integration, the CNR performance can be overall improved in terms of communication coverage, positioning resolution, and remote sensing response time, etc. A typical potential of such integration is to leverage the broadband communication capability to perform resolution enhanced multi-satellite sensing fusion and low-latency transportation sensing information downloading for automated driving applications.
  • L2: Signal-Level Integration
In this level, the signals for the CNR are expected to be combined, i.e., the signal emitted by the satellites carries communication information, navigation reference, and remote sensing probes simultaneously, which leads to a more efficient use of limited spectrum resource than the function integration. We can also have an easier variant that flexibly reuses the wireless resource in time, frequency, or space domains. Generally, there is a fundamental tradeoff among different performance of the CNR in designing multi-purpose signals. In practical systems, how functionalities are combined in the signal should depend on the realtime geographical distribution of the CNR requirements. Particularly, for a time instance at which only communication and navigation services are required by the ground users in a region, there is no need to embrace the remote sensing function in the current signal design. The vehicular terminals can benefit greatly from equivalent enlarged resource pool with signal-level integration, which holds attractive potentials to provide high-resolution traffic sensing information and lane-level navigation.
  • L3: Data-level Integration
Different from the above two levels, data-level integration allows the satellites to share information collected by different functionalities, where the information coming originally from one functionality may be very helpful in optimizing the network performance from the perspective of another functionality. The remote sensing results generally show the geographical vehicular traffic distribution, which can be leveraged to conduct more efficient traffic-aware communication and navigation resource scheduling. Using navigation information, such as location and speed of vehicles, the communication channel may be more efficiently predicted, which allows more adaptive communication transceiver design. By analyzing the communication channel state information (CSI) of a large number of vehicular users, we can estimate the overall channel conditions for a given region, which can be exploited to perform better remote sensing beamforming design and CSI fingerprinting based positioning. Finally, high mobility of ground vehicles causes rapid change of remote transportation sensing images, navigation GDOP values, and communication load distribution, which gives rise to a high demand on fast inter-functionality data mining.
  • L4: Application-Level Integration
All information delivery, localization, and earth surface observation functionalities are actually supporting technologies rather than the final objectives of the ICNR system. To this end, the application-level integration aims to achieve the ultimate goal on providing satisfied services for various applications of ground users. As previously mentioned, an autonomous driving application simultaneously requires the CNR services. Therefore, the application-level integration takes the waveform design, resource allocation, and data exploitation as a whole to jointly optimize all procedures, as illustrated in Fig. 3. Particularly, coordination and synchronization among different functionalities within each satellite and even among different satellites are particularly needed to avoid interference and conflict. In such cases, providing on-demand services is one of the key features of the ICNR networks, implying that an increasing demand on one of the three functionalities triggers the system resource allocation tilting to it. As future vehicular applications often require comprehensive CNR services for autonomous driving, high reliability and low latency are usually required in communication, positioning, and remote transportation sensing. The goal-oriented application-level integration holds great potentials to jointly satisfy such heterogeneous requirements by targeted resource scheduling.
TABLE I: Network Setup of the Case Study
Requirement/Parameter Value or Description
[$M_{c},M_{n},M_{r}$] $[4500,250,250]$
[$B_{c},B_{n},B_{r}$] $[250,25,25]$ MHz
Communication transmit power $20$ W
Navigation transmit power $20$ W
Remote sensing transmit power $80$ W
Navigation ranging error of unit distance $2.28\times 10^{-8}$ m
Communication elevation angle $60^{\circ}$
Navigation elevation angle $20^{\circ}$
Remote sensing view angle $30^{\circ}$
Remote sensing swath width $20$ km
Communication minimum required rate $0.7$ Mbps
Remote sensing data size 360 Gbits
Remote sensing data delivery rate 1 bps per Hz
Large-scale fading $110+37.6\log_{10}(d)$
Small-scale fading Rayleigh fading
IV A Fundamental Case Study
This section evaluates performance gain achieved by the lowest two levels of integration. For the traditional situation without integration, denote the numbers of the CNR satellites to be $M_{c}$, $M_{n}$, and $M_{r}$, respectively, and the available spectra for the CNR satellites to be $B_{c}$, $B_{n}$, and $B_{r}$, respectively. Each type of the satellites is uniformly distributed on the sphere with an altitude of 500 km. In the case of function-level integration, there are equivalent $M_{c}+M_{n}+M_{r}$ satellites uniformly distributed on the same sphere, each of which has the CNR functionalities simultaneously. One step further, we assume ideal waveform design such that all the CNR functionalities can share the $B_{c}+B_{n}+B_{r}$ bandwidth without performance tradeoffs in the signal-level integration. There are 1,400 vehicular users distributed on circular cap area of $7.36\times 10^{5}\ {\rm km}^{2}$ near the equator. Each user has a minimum required data rate for communication and a data rate lower than this threshold causes service outage. Meanwhile, every user needs positioning, timing, and remote sensing information downloading, where the LEO enhanced positioning is performed on the basis of the GPS. In the integration cases, we consider each swath is illuminated by two satellites simultaneously to get observation resolution improvement, say, by a half, in the range direction by data fusion. The AoI of the remote sensing data received by the ground users is derived by adding satellite revisit time and the data downloading time. Some key simulation parameters are summarized in Table I.
Figure 4: Estimated gain of function-level and signal-level integration.
The overall performance of the CNR is presented by a spider web shown in Fig. 4, where the communication is evaluated by capacity outage probability and mean ergodic capacity marked in red, the navigation is characterized by positioning accuracy and timing error marked in blue, and the remote sensing is captured by resolution in range direction and AoI measured at the ground receivers marked in purple. It can be observed that both function-level and signal-level integration can comprehensively improve the CNR performance. The function-level integration introduces more available satellites for the three functionalities, which shortens the communications distance, provides more anchor location references for navigation, and decreases the revisit time for remote sensing. The gain of the signal-level integration mainly comes from expansion of the available spectrum for each functionality, which results in less improvement in communication regime as the increased amount of spectrum is smaller than other two functionalities. More importantly, the almost centimeter-level positioning achieved by the signal-level integration is quite fascinating for autonomous vehicles to perform lane navigation. Meanwhile, the decimeter-level earth observation accuracy is helpful to distinguish different vehicles on the ground, which can be leveraged to conduct sophisticated road traffic analysis and transportation system optimization. Such quantitative improvement of the navigation and remote sensing performance by even the lowest two levels of integration, e.g., decreasing the positioning error and view resolution by one order of magnitude, shines light on the future ICNR industry while communication performance is also expected to be enhanced significantly by deeply exploiting the navigation and remote sensing data in higher levels of integration. Finally, it is safe to admit the promising potentials of the ICNR in comprehensively enhancing the CNR performance as the integration level goes up, especially when the total number of the LEO satellites grows further to tens of thousands, which is promising to offer distinct transportation observation, centimeter-level positioning, and timely downloading of huge remote sensing data for future vehicular applications.
V Key Tradeoffs of ICNR Management
Optimization of the ICNR system has to jointly take the three functionalities into account. However, due to distinct working mechanisms of the three functionalities from one another, many performance tradeoffs need to be addressed, especially for vehicular applications.
V-A Constellation Design
As the constellation size or orbit altitude increases, the communication coverage probability usually increases first due to reducing coverage holes and then decreases due to rising interference if no advanced interference coordination technique is implemented. On the other hand, the GDOP in the navigation regime monotonically decreases with the constellation size and the orbit altitude, and around 600 satellites deployed in the orbits with an altitude of 500 km can achieve the same performance with the GPS system [11]. However, only linear decrease of the GDOP may be achieved by exponential increase of the number of satellites greater than about 600. Furthermore, in addition to fast varying satellite distribution, the mobility of vehicles makes terminals experience more complicated positioning performance influenced by the constellation. For the remote sensing functionality, the constellation size has a direct influence on the temporal resolution in terms of the revisit time but not on the spatial resolution, where the revisit time in strip-map mode is generally discretely decreased with the number of satellites. It is worth noting that on-demand spotlight mode may be more suitable for transportation remote sensing for vehicular applications, where urban area has a higher requirement than rural area. Since a mega-constellation of LEO satellites must be launched batch by batch lasting for several years or more, it is thus critical to optimize the evolution route of the error-growing constellation, so as to march in lockstep with the increasing demand on the CNR in autonomous driving.
V-B Satellite Association
Associating a terminal to the nearest access point has long been the fundamental principle in wireless communication field to experience the lowest channel fading, implying that the nadir of a satellite on the ground gains the best channel quality. However, the thing becomes totally different for the navigation and remote sensing. According to (4), directly looking below with a view angle of $\varphi=0$ fails to distinguish any two objects with any distance apart. Therefore, a SAR satellite needs to take a side look, and a typical view angle of $\varphi=30^{\circ}$ at an orbit altitude of $800$ km could lead to a distance of around 373 km between the swath center and the nadir [11]. Although the ground range resolution gets better as the view angle becomes larger, a moderate view angle is usually adopted to avoid severe signal attenuation. As for the navigation, the GDOP becomes better if the satellites are more scattered throughout the sky, indicating that the navigation functionality prefers to use signals emitted from far satellites dispersed in all directions. In other words, the navigation terminals may need to discard some signals from nearby satellites although many satellites can be seen in the era of mega-constellations of LEO satellites. The distinct association preference among different functionalities, as shown in Fig. 5, causes great difficulty in multi-purpose signal design and beam control toward moving vehicles. Additionally, on the terminal side, vehicles need to dynamically conduct satellite selection for each functionally to adapt to satellite distribution and own location variations.
Figure 5: Preferable satellite association relationship for CNR functionalities.
V-C Resource Management
On-demand service providing could probably be one of the main features of the LEO mega-constellation based ICNR systems. Particularly, as a small number of satellites are deployed in traditional remote sensing, the strip-map mode is commonly implemented to obtain earth observation images with global coverage, usually causing long revisit time greater than a couple of days or more. By contrast, since every ground point is able to see many satellites in the LEO mega-constellations, the spotlight mode is more appropriate to be performed whenever there is a particular demand on real-time observation towards a region while the strip-map mode is still conducted by a number of satellites to offer regular earth observation. Therefore, how many satellites work in the strip-map mode and how many are reserved for sporadic services should be carefully designed. In addition, as the CNR functionalities share the limited resources, such as power, computing, cache, and spectrum, smart allocation of these heterogeneous resources plays a prominent role in enhancing overall performance, including how to balance communication and computing payload by appropriately processing the remote sensing data in orbit [15]. Since a growing number of intelligent vehicles persistently impose requirements on high-quality earth observation, navigation, and remote sensing data downloading, sequential resource management decisions have to be optimized in response to network environment variations, which is pretty challenging in the highly dynamic ICNR networks with fast moving satellites and vehicles.
VI Conclusion
In this article, we have proposed the ICNR framework in LEO satellite networks with vehicular applications. After presenting the most fundamental principles of the CNR functionalities, we have classified the proposed ICNR networks in different integration levels. The function-level and signal-level integration have then been instantiated as a numerical case study to show the integration gain. Finally, several key tradeoffs in managing the ICNR networks have been emphasized from the perspective of vehicular applications.